OpenStudy (anonymous):

Solve for a and b using properties of 45-45-90 triangles

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

and you need to know again that sin45=sqrt2/2 using this will get a= b*sqrt2/2 and b= c/sqrt2 /2 = 2c/sqrt2